OpenStudy (anonymous):

find all solutions to the equation 2cos(x) -1 =0 on the interval [-4pi,-2pi]?

OpenStudy (zzr0ck3r):

2cos(x)-1=0 cos(x)=1/2 this is true for x=pi/3 +- 2pi and this is true for (5/3)pi+-2pi

OpenStudy (zzr0ck3r):

so take pi/3 and subtract 2pi and take all the values in that interval do the same for (5/3)pi